two brothers boll and eric are 4 years apart and bill is the oldest one if the same of the boys age is 28 how old is bill and eric?
Answer by ewatrrr(24785)


Hi
bill and eric are 4 years apart
Let x and (x+4)represent Eric's and Bill's ages respectively
Question states***
x + (x+4)= 8
Solving for x
2x + 4 = 28
2x = 24
x = 12,Eric's age. Bill is 16 (12+4)
CHECKING our Answer***
12 + 16 = 28